# The Jordan canonical form of the Fréchet derivative of a matrix function

### v1

@@ Abstract
Let $\mathbb{F}$ be an algebraically closed field of characteristic 0. Given a square matrix $A\in\mathbb{F}^{n\times n}$ and a polynomial $f\in\mathbb{F}[w]$, we determine the Jordan canonical form of the formal Fréchet derivative of $f(A)$, in terms of that of $A$ and of $f$. When $\mathbb{F}\subseteq\mathbb{C}$, via Hermite interpolation, our result provides a full solution to [N.J. Higham, Functions of Matrices: Theory and Computation, Research Problem 3.11]. A generalization consists of finding the Jordan canonical form of linear combinations of Kronecker products of powers of two square matrices, i.e., $\sum_{i,j}a_{ij}(X^{i}\otimes Y^{j})$. For this generalization, we provide some new partial results, including a partial solution under certain assumptions and general bounds on the number and the sizes of Jordan blocks.
@@ 1 Motivation and statement of the problem
The operator $Df_{w}$, when it exists, is unique and is known as the Fréchet derivative of $f$ at $w$. When $U=V=\mathbb{C}^{n\times n}$, a case of particular interest is when $f$ is a matrix function in the linear algebraic sense, which is a more restrictive notion than just applying the usual definition of a function to a matrix [5, 7]. Since $U$ and $V$ are now normed vector spaces of finite dimension $n^{2}$ over $\mathbb{C}$, it is possible to represent the linear map $E\rightarrow Df_{W}(E)$ as an $n^{2}\times n^{2}$ matrix. Clearly, the representation depends on the choice of a basis for $\mathbb{C}^{n\times n}$. The most common choice is induced by the canonical basis of $\mathbb{C}^{n}$ via the two formal operations of Kronecker product and vec operator [5, p. 331]. Such representation of $Df_{W}$ is sometimes referred to as the Kronecker form of the Fréchet derivative. Furthermore, if $f$ is regular enough, say, analytic in an open set including the eigenvalues of $W$, then there exists a polynomial $p$ (depending on both $f$ and $W$) such that $Df_{W}=Dp_{W}$ [7, Theorem 6.6.14]; see also the discussion in [5, Section 3.2]. Therefore studying the case of polynomial functions suffices to cover a vast class of functions. It can be shown [5, Problem 3.6] that for the polynomial function $f(W)=\sum_{i=-1}^{D-1}f_{i}W^{i+1}$ the Kronecker form of $Df_{W}$ is the matrix
@@
We are especially interested in describing the Jordan canonical form (JCF) of (3) in terms of the scalar polynomial (2) and the JCFs of $X$ and $Y$. It is clear that $p(\lambda,\mu)$ is an eigenvalue of (3) if and only if $\lambda$ is an eigenvalue of $X$ and $\mu$ is an eigenvalue of $Y$. Hence, the only nontrivial part of the computation of the JCF of (3) is to determine the sizes of each Jordan block. We call this problem the bivariate Jordan problem. Some specific instances of the bivariate Jordan problem have been solved long ago. If the only nonzero coefficient in (3) is $a_{11}=1$, we obtain the Kronecker product of two matrices, whose JCF has been thoroughly studied by various authors [1, 3, 8, 10, 12], and for this case the solution is well known. Another important special case is the Kronecker sum of two matrices: $a_{10}=a_{01}=1$, $a_{ij}=0$ otherwise. Kronecker sums are often used and a complete characterization of their JCF first appeared in [12, Theorem 2].
@@
$P(X,Y)=\sum_{i=0}^{D}f_{i}\left(\sum_{j=0}^{i}X^{i}\otimes Y^{j}\right)\ \in\mathbb{F}^{mn\times mn}.$ (4)
@@
we also call Fréchet-Jordan problem the special case of the bivariate Jordan problem where (3) takes the form (4). Our motivation is to answer [5, Research Problem $3.11$], in which N. Higham asked this question for the Kronecker form of the Fréchet derivative of a matrix function. In other words, [5, Research Problem $3.11$] is the Fréchet-Jordan problem for $\mathbb{F}=\mathbb{C}$ and $X=Y^{T}$. To our knowledge, [5, Research Problem $3.11$] was until now still an open question. The present paper fully solves it.
@@
    For the special case (4), we give a full general solution to the Fréchet-Jordan problem, regardless of whether (6) holds or fails. By taking $\mathbb{F}=\mathbb{C}$ and $X=Y^{T}$ in (4), this in particular solves [5, Research Problem 3.11].
We structure our exposition as follows. We make some preliminary analysis of the bivariate Jordan problem in Section 2. In Subsection 3.1 we solve the bivariate Jordan problem assuming that (6) holds for all eigenvalue pairs ($\lambda,\mu)$, which as discussed happens for a non-empty and Zariski-open subset of problem instances of bounded size and degree. In Subsection 3.2 we give general bounds for the degenerate case when the assumption is not true. In Section 4, we fully solve the Fréchet-Jordan problem. Finally, some conclusions are drawn in Section 5.
@@ 2 Simplifying preliminaries
Suppose that $X=SJ_{X}S^{-1}$ and $Y=RJ_{Y}R^{-1}$ for certain nonsingular matrices $S$ and $R$, with $J_{X},J_{Y}$ both in JCF. Then by [7, Lemma 4.2.10] $(S^{-1}\otimes R^{-1})(X^{i}\otimes Y^{j})(S\otimes R)=J_{X}^{i}\otimes J_{Y}^{j}$. Therefore, $P(X,Y)$ and $P(J_{X},J_{Y})$ are similar, which means that there is no loss of generality in taking both $X$ and $Y$ in JCF. Furthermore, it suffices to separately analyze each pair $(X_{i},Y_{j})$ where $X_{i}$ (resp. $Y_{j}$) is a Jordan block of $X$ (resp. $Y$).
@@
[7, Corollary 4.3.16] Suppose that $X=\bigoplus_{i}X_{i}$ and $Y=\bigoplus_{j}Y_{j}$ are block diagonal. Then $X\otimes Y$ is similar to $\bigoplus_{i,j}(X_{i}\otimes Y_{j})$. Moreover, the similarity matrix is a permutation matrix that only depends on the sizes of the blocks $X_{i}$ and $Y_{j}$.
@@
Here and below, $N_{k}=J_{k}(0)$ denotes a nilpotent Jordan block of size $k$. Denoting formal partial derivatives with resepct to $x$ by $\frac{\partial}{\partial x}$, it can be checked by direct computation that $J_{m}(\lambda)^{i}$ is the upper triangular Toeplitz matrix whose first row is $\begin{bmatrix}x^{i}&\frac{\partial x^{i}}{\partial x}&\frac{1}{2!}\frac{\partial^{2}x^{i}}{\partial^{2}x}&\dots&\frac{1}{i!}\frac{\partial^{m}x^{i}}{\partial^{m}x}\end{bmatrix}$ evaluated at $x=\lambda$; an analogous result holds for $J_{n}(\mu)^{j}$. By the definition of the Kronecker product one has that $J_{m}(\lambda)^{i}\otimes J_{n}(\mu)^{j}$ is equal to the evaluation at $x=\lambda$ of the matrix
@@
Proof. Recall [7, Corollary 4.3.10] that, for any two square matrices $X$ and $Y$, $X\otimes Y$ and $Y\otimes X$ are permutation similar and the similarity matrix $\Pi$ only depends on the sizes of $X$ and $Y$. Comparing (2) with $q(x,y)=\sum_{0\leq i,j\leq k}a_{ji}x^{i}y^{j}$, we get that $Q(Y,X)=\sum_{0\leq i,j\leq k}a_{ji}Y^{i}\otimes X^{j}=\sum_{0\leq i,j\leq k}a_{ij}Y^{j}\otimes X^{i}=\Pi^{T}P(X,Y)\Pi$. The explicit form of $\Pi$ is given in [7, Equation 4.3.9b]: It is straightforward to check that the formula given there is equivalent to the one given here.
@@
[6, Section 0.9.7] The set of $n\times n$ upper triangular Toeplitz matrices with elements in a commutative ring, equipped with matrix addition and matrix multiplication, is itself a commutative ring.
@@
There is no loss of generality in assuming $A_{0}=0$. Indeed, we may equivalently solve $(Z_{m-1}-I_{m}\otimes A_{0})X=X(Z_{1}-I_{m}\otimes A_{0})$ because every matrix in $R^{m\times m}$ commutes with $I_{m}\otimes A_{0}$. Let $U=N_{m}\otimes I_{n}\in\hat{R}$ and, for $k=1,\dots,m-1$, define $Z_{k}=\sum_{i=1}^{k}A_{i}U^{i}$. (Note that this definition also covers $Z_{1},Z_{m-1}$.) The action of $U$ when acting from the left is an upwards row shift, while the action of $U$ from the right is a rightwards columnn shift. Since $R$ is a commutative ring we have [^1] Here and below the notation $X(a:b,c:d)$, where $a\leq b,c\leq d$ are positive integers, indicates the submatrix of $X$ obtained by selecting the rows with indices $a,\dots,b$ and the columns with indices $c,\dots,d$. $XZ_{1}=A_{1}\begin{bmatrix}0&X(1:m-1,1:m-1)\\ 0&0\end{bmatrix}$. On the other hand, $Z_{m-1}X=\sum_{i=1}^{m-1}A_{i}\begin{bmatrix}0&X(i+1:m,i+1:m)\\ 0&0\end{bmatrix}$ ( the blocks have different sizes for different terms in the summation). Therefore, $D_{ij}=0$ if ji<2j-i<2, while $D_{ij}=A_{j-i}-A_{1}X_{i,j-1}+\sum_{\ell=1}^{j-i-1}(A_{\ell}X_{i+\ell,j})$ if $j-i\geq 2$. Thus, $D_{ij}$ is an affine function $A_{1}$: $D_{ij}=D_{ij}^{(0)}+D_{ij}^{(1)}A_{1}$. In particular, when $j\geq i+2$, $D_{ij}^{(1)}=X_{i+1,j}-X_{i,j-1}$ and $D_{i,j}^{(0)}=\sum_{\ell=2}^{j-i-1}(A_{\ell}X_{i+\ell,j})$. Both $X_{i+1,j}$ and $X_{i,j-1}$ belong to the (ji1)(j-i-1)th superdiagonal [^2] The first superdiagonal of a square matrix is the set of the elements directly above the diagonal. Recursively, the $k$th superdiagonal is defined as the set of the elements directly above the (k1)(k-1)th superdiagonal. of $X$. Moreover, the unknown elements of $X$ appearing in $D^{(0)}_{ij}$ must belong to one of the first ji2j-i-2 superdiagonals of $X$. We conclude that the following procedure can be applied to solve the problem:
@@ 3.1 Generic case: Full solution
The first item in Theorem 3.1, i.e., the case $p_{x}\neq 0\neq p_{y}$, is not a new result. It is indeed equivalent to [11, Corollary 1].
@@
Consider, as in [5, Research Problem 3.11], the Fréchet derivative of $f(W)=W^{2}$ for $W=J_{2}(0)=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$. The associated bivariate polynomial is $p(x,y)=x+y$, therefore $p_{x}=p_{y}=1$. Since there is a single Jordan block of size $2$ and eigenvalue 0, Theorem 3.1 correctly predicts that the JCF of (1) is $J_{3}(0)+J_{1}(0)$.
@@
[12, Theorem 2] For each Jordan block of $X$ of size $n$ with eigenvalue $\lambda$ and for each Jordan block of $Y$ of size $m$ with eigenvalue $\mu$, the JCF of the Kronecker sum $X\otimes I+I\otimes Y$ there are, associated with each eigenvalue $\lambda+\mu$, $\min(m,n)$ Jordan blocks of sizes $m+n+1-2\min(m,n),m+n+3-2\min(m,n),\dots,m+n-1$.
@@ 3.1.1 The Jordan canonical form of $P(J_{m}(\lambda),J_{n}(\mu))$ for $p_{x}\neq 0$, $p_{y}\neq 0$
As discussed, this case appeared as [11, Corollary 1]. To make the paper self-contained, we nevertheless give an independent proof. Since $p_{x}\neq 0$, we can apply Lemma 2.4 and claim that it suffices to find the JCF of $I_{m}\otimes B+N_{m}\otimes I_{n}$, where $B$ is defined in (5). Then, since $p_{y}\neq 0$, we can apply Lemma 2.4 again to reduce $B$ to bidiagonal form via a certain similarity matrix $X$. Hence, by applying the similarity transformation associated with $I_{m}\otimes X$ to our larger matrix, we reduce to the problem of finding the JCF of
@@
The matrix above is a Kronecker sum, and it is known that the lengths of its Jordan chains are m+n+12km+n+1-2k for $k=1,2,\dots,\min(m,n)$. This is stated in [12, Theorem 2] (a crucial claim in [12] is verified in [10]). An alternative approach is discussed in [3], where it is shown that this problem is related to the study of the $k$-paths of the Cartesian product of $\Gamma_{m}$ and $\Gamma_{n}$, where $\Gamma_{k}$ is the digraph having vertices $1,2,\dots,k$ and arcs $12,23,\dots,(k-1)k$. We refer the reader to [3] for definitions and details. A slight modification to the proof [3, Lemma 4.3] leads to the same conclusion as in [12, Theorem 2]. We also note that, substituting every occurrence of $I_{n}$ in the first block superdiagonal with $I_{n}+N_{n}$ via a block-diagonal similarity, we could equivalently have used the results concerning Kronecker products: see [3, Lemma 4.3], and also [1, 8, 10, 12]. Therefore the first item in Theorem 3.1 is proved.

### v2

@@ Abstract
Let $\mathbb{F}$ be an algebraically closed field of characteristic 0. Given a square matrix $A\in\mathbb{F}^{n\times n}$ and a polynomial $f\in\mathbb{F}[w]$, we determine the Jordan canonical form of the formal Fréchet derivative of $f(A)$, in terms of that of $A$ and of $f$. When $\mathbb{F}\subseteq\mathbb{C}$, via Hermite interpolation, our result provides a solution to [N.J. Higham, Functions of Matrices: Theory and Computation, Research Problem 3.11]. A generalization consists of finding the Jordan canonical form of linear combinations of Kronecker products of powers of two square matrices, i.e., $\sum_{i,j}a_{ij}(X^{i}\otimes Y^{j})$. For this generalization, we provide some new partial results, including a partial solution under certain assumptions and general bounds on the number and the sizes of Jordan blocks.
@@ 1 Motivation and statement of the problem
The operator $Df_{w}$, when it exists, is unique and is known as the Fréchet derivative of $f$ at $w$. When $U=V=\mathbb{C}^{n\times n}$, a case of particular interest is when $f$ is a matrix function in the linear algebraic sense, which is a more restrictive notion than just applying the usual definition of a function to a matrix [6, 8]. Since $U$ and $V$ are now normed vector spaces of finite dimension $n^{2}$ over $\mathbb{C}$, it is possible to represent the linear map $E\rightarrow Df_{W}(E)$ as an $n^{2}\times n^{2}$ matrix. Clearly, the representation depends on the choice of a basis for $\mathbb{C}^{n\times n}$. The most common choice is induced by the canonical basis of $\mathbb{C}^{n}$ via the two formal operations of Kronecker product and vec operator [6, p. 331]. Such representation of $Df_{W}$ is sometimes referred to as the Kronecker form of the Fréchet derivative. Furthermore, if $f$ is regular enough, say, analytic in an open set including the eigenvalues of $W$, then there exists a polynomial $p$ (depending on both $f$ and $W$) such that $Df_{W}=Dp_{W}$ [8, Theorem 6.6.14]; see also the discussion in [6, Section 3.2]. Therefore studying the case of polynomial functions suffices to cover a vast class of functions. It can be shown [6, Problem 3.6] that for the polynomial function $f(W)=\sum_{i=-1}^{D}f_{i}W^{i+1}$ the Kronecker form of $Df_{W}$ is the matrix
@@
We are especially interested in describing the Jordan canonical form (JCF) of (3) in terms of the scalar polynomial (2) and the JCFs of $X$ and $Y$. It is clear that $p(\lambda,\mu)$ is an eigenvalue of (3) if and only if $\lambda$ is an eigenvalue of $X$ and $\mu$ is an eigenvalue of $Y$. Hence, the only nontrivial part of the computation of the JCF of (3) is to determine the sizes of each Jordan block. We call this problem the bivariate Jordan problem. Some specific instances of the bivariate Jordan problem have been solved long ago. If the only nonzero coefficient in (3) is $a_{11}=1$, we obtain the Kronecker product of two matrices, whose JCF has been thoroughly studied by various authors [1, 4, 9, 11, 13], and for this case the solution is well known. Another important special case is the Kronecker sum of two matrices: $a_{10}=a_{01}=1$, $a_{ij}=0$ otherwise. Kronecker sums are often used and a complete characterization of their JCF first appeared in [13, Theorem 2].
@@
$P(X,Y)=\sum_{i=0}^{D}f_{i}\left(\sum_{j=0}^{i}X^{j}\otimes Y^{i-j}\right)\ \in\mathbb{F}^{mn\times mn}.$ (4)
@@
we also call Fréchet-Jordan problem the special case of the bivariate Jordan problem where (3) takes the form (4). Our motivation is to answer [6, Research Problem $3.11$], in which N. Higham asked this question for the Kronecker form of the Fréchet derivative of a matrix function. In other words, [6, Research Problem $3.11$] is the Fréchet-Jordan problem for $\mathbb{F}=\mathbb{C}$ and $X=Y^{T}$. To our knowledge, [6, Research Problem $3.11$] was until now still an open question. The present paper answers it.
@@
    For the special case (4), we give a full general solution to the Fréchet-Jordan problem, regardless of whether (6) holds or fails. Our solution is by a closed formula when $X$ and $Y$ do not share eigenvalues, whereas for shared eigenvalues it reduces the problem to computing the ranks of certain banded integer Toeplitz matrices (of much smaller size than (4)). By taking $\mathbb{F}=\mathbb{C}$ and $X=Y^{T}$ in (4), this in particular solves [6, Research Problem 3.11].
We structure our exposition as follows. We make some preliminary analysis of the bivariate Jordan problem in Section 2. In Subsection 3.1 we solve the bivariate Jordan problem assuming that (6) holds for all eigenvalue pairs ($\lambda,\mu)$, which as discussed happens for a non-empty and Zariski-open subset of problem instances of bounded size and degree. In Subsection 3.2 we give general bounds for the degenerate case when the assumption is not true. In Section 4, we fully solve the Fréchet-Jordan problem in the sense mentioned above. Finally, some conclusions are drawn in Section 5.
@@ 2 Simplifying preliminaries
Suppose that $X=SJ_{X}S^{-1}$ and $Y=RJ_{Y}R^{-1}$ for certain nonsingular matrices $S$ and $R$, with $J_{X},J_{Y}$ both in JCF. Then by [8, Lemma 4.2.10] $(S^{-1}\otimes R^{-1})(X^{i}\otimes Y^{j})(S\otimes R)=J_{X}^{i}\otimes J_{Y}^{j}$. Therefore, $P(X,Y)$ and $P(J_{X},J_{Y})$ are similar, which means that there is no loss of generality in taking both $X$ and $Y$ in JCF. Furthermore, it suffices to separately analyze each pair $(X_{i},Y_{j})$ where $X_{i}$ (resp. $Y_{j}$) is a Jordan block of $X$ (resp. $Y$).
@@
[8, Corollary 4.3.16] Suppose that $X=\bigoplus_{i}X_{i}$ and $Y=\bigoplus_{j}Y_{j}$ are block diagonal. Then $X\otimes Y$ is similar to $\bigoplus_{i,j}(X_{i}\otimes Y_{j})$. Moreover, the similarity matrix is a permutation matrix that only depends on the sizes of the blocks $X_{i}$ and $Y_{j}$.
@@
Here and below, $N_{k}=J_{k}(0)$ denotes a nilpotent Jordan block of size $k$. Denoting formal partial derivatives with respect to $x$ by $\frac{\partial}{\partial x}$, it can be checked by direct computation that $J_{m}(\lambda)^{i}$ is the upper triangular Toeplitz matrix whose first row is $\begin{bmatrix}x^{i}&\frac{\partial x^{i}}{\partial x}&\frac{1}{2!}\frac{\partial^{2}x^{i}}{\partial^{2}x}&\dots&\frac{1}{i!}\frac{\partial^{m}x^{i}}{\partial^{m}x}\end{bmatrix}$ evaluated at $x=\lambda$; an analogous result holds for $J_{n}(\mu)^{j}$. By the definition of the Kronecker product one has that $J_{m}(\lambda)^{i}\otimes J_{n}(\mu)^{j}$ is equal to the evaluation at $x=\lambda$ of the matrix
@@
Proof. Recall [8, Corollary 4.3.10] that, for any two square matrices $X$ and $Y$, $X\otimes Y$ and $Y\otimes X$ are permutation similar and the similarity matrix $\Pi$ only depends on the sizes of $X$ and $Y$. Comparing (2) with $q(x,y)=\sum_{0\leq i,j\leq k}a_{ji}x^{i}y^{j}$, we get that $Q(Y,X)=\sum_{0\leq i,j\leq k}a_{ji}Y^{i}\otimes X^{j}=\sum_{0\leq i,j\leq k}a_{ij}Y^{j}\otimes X^{i}=\Pi^{T}P(X,Y)\Pi$. The explicit form of $\Pi$ is given in [8, Equation 4.3.9b]: It is straightforward to check that the formula given there is equivalent to the one given here.
@@
[7, Section 0.9.7] The set of $n\times n$ upper triangular Toeplitz matrices with elements in a commutative ring, equipped with matrix addition and matrix multiplication, is itself a commutative ring.
@@
There is no loss of generality in assuming $A_{0}=0$. Indeed, we may equivalently solve $(Z_{m-1}-I_{m}\otimes A_{0})X=X(Z_{1}-I_{m}\otimes A_{0})$ because every matrix in $R^{m\times m}$ commutes with $I_{m}\otimes A_{0}$. Let $U=N_{m}\otimes I_{n}\in\hat{R}$ and, for $k=1,\dots,m-1$, define $Z_{k}=\sum_{i=1}^{k}A_{i}U^{i}$. (Note that this definition also covers $Z_{1},Z_{m-1}$.) The action of $U$ when acting from the left is an upwards row shift, while the action of $U$ from the right is a rightwards column shift. Since $R$ is a commutative ring we have [^1] Here and below the notation $X(a:b,c:d)$, where $a\leq b,c\leq d$ are positive integers, indicates the submatrix of $X$ obtained by selecting the rows with indices $a,\dots,b$ and the columns with indices $c,\dots,d$. $XZ_{1}=A_{1}\begin{bmatrix}0&X(1:m-1,1:m-1)\\ 0&0\end{bmatrix}$. On the other hand, $Z_{m-1}X=\sum_{i=1}^{m-1}A_{i}\begin{bmatrix}0&X(i+1:m,i+1:m)\\ 0&0\end{bmatrix}$ ( the blocks have different sizes for different terms in the summation). Therefore, $D_{ij}=0$ if ji<2j-i<2, while $D_{ij}=A_{j-i}-A_{1}X_{i,j-1}+\sum_{\ell=1}^{j-i-1}(A_{\ell}X_{i+\ell,j})$ if $j-i\geq 2$. Thus, $D_{ij}$ is an affine function $A_{1}$: $D_{ij}=D_{ij}^{(0)}+D_{ij}^{(1)}A_{1}$. In particular, when $j\geq i+2$, $D_{ij}^{(1)}=X_{i+1,j}-X_{i,j-1}$ and $D_{i,j}^{(0)}=\sum_{\ell=2}^{j-i-1}(A_{\ell}X_{i+\ell,j})$. Both $X_{i+1,j}$ and $X_{i,j-1}$ belong to the (ji1)(j-i-1)th superdiagonal [^2] The first superdiagonal of a square matrix is the set of the elements directly above the diagonal. Recursively, the $k$th superdiagonal is defined as the set of the elements directly above the (k1)(k-1)th superdiagonal. of $X$. Moreover, the unknown elements of $X$ appearing in $D^{(0)}_{ij}$ must belong to one of the first ji2j-i-2 superdiagonals of $X$. We conclude that the following procedure can be applied to solve the problem:
@@ 3.1 Generic case: Full solution
The first item in Theorem 3.1, i.e., the case $p_{x}\neq 0\neq p_{y}$, is not a new result. It is indeed equivalent to [12, Corollary 1].
@@
Consider, as in [6, Research Problem 3.11], the Fréchet derivative of $f(W)=W^{2}$ for $W=J_{2}(0)=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}$. The associated bivariate polynomial is $p(x,y)=x+y$, therefore $p_{x}=p_{y}=1$. Since there is a single Jordan block of size $2$ and eigenvalue 0, Theorem 3.1 correctly predicts that the JCF of (1) is $J_{3}(0)+J_{1}(0)$.
@@
[13, Theorem 2] For each Jordan block of $X$ of size $n$ with eigenvalue $\lambda$ and for each Jordan block of $Y$ of size $m$ with eigenvalue $\mu$, the JCF of the Kronecker sum $X\otimes I+I\otimes Y$ there are, associated with each eigenvalue $\lambda+\mu$, $\min(m,n)$ Jordan blocks of sizes $m+n+1-2\min(m,n),m+n+3-2\min(m,n),\dots,m+n-1$.
@@ 3.1.1 The Jordan canonical form of $P(J_{m}(\lambda),J_{n}(\mu))$ for $p_{x}\neq 0$, $p_{y}\neq 0$
As discussed, this case appeared as [12, Corollary 1]. To make the paper self-contained, we nevertheless give an independent proof. Since $p_{x}\neq 0$, we can apply Lemma 2.4 and claim that it suffices to find the JCF of $I_{m}\otimes B+N_{m}\otimes I_{n}$, where $B$ is defined in (5). Then, since $p_{y}\neq 0$, we can apply Lemma 2.4 again to reduce $B$ to bidiagonal form via a certain similarity matrix $X$. Hence, by applying the similarity transformation associated with $I_{m}\otimes X$ to our larger matrix, we reduce to the problem of finding the JCF of
@@
The matrix above is a Kronecker sum, and it is known that the lengths of its Jordan chains are m+n+12km+n+1-2k for $k=1,2,\dots,\min(m,n)$. This is stated in [13, Theorem 2] (a crucial claim in [13] is verified in [11]). An alternative approach is discussed in [4], where it is shown that this problem is related to the study of the $k$-paths of the Cartesian product of $\Gamma_{m}$ and $\Gamma_{n}$, where $\Gamma_{k}$ is the digraph having vertices $1,2,\dots,k$ and arcs $12,23,\dots,(k-1)k$. We refer the reader to [4] for definitions and details. A slight modification to the proof [4, Lemma 4.3] leads to the same conclusion as in [13, Theorem 2]. We also note that, substituting every occurrence of $I_{n}$ in the first block superdiagonal with $I_{n}+N_{n}$ via a block-diagonal similarity, we could equivalently have used the results concerning Kronecker products: see [4, Lemma 4.3], and also [1, 9, 11, 13]. Therefore the first item in Theorem 3.1 is proved.